OpenStudy (anonymous):

Having trouble using IS\OF = X\100 to set up this problem.... The Community of Bardville has 1332 Left Handed Females, if 48% of the community is Female, and 15% of all females are left handed, how many people are in the community?

OpenStudy (triciaal):

let x = number of females 15 % of the females are left-handed = 1332 15/100 * x = 1332 x = 8880 let the number of people in the community = y 48/100 = 8880/y y = 8880*100/48 = 18500 another approach 15% of 48% of number of people = 1332 number of people = 1332/ 15%*48% = 1332/0.072 number of people in the community = 18500

OpenStudy (anonymous):

Super helpful thank you!!

OpenStudy (triciaal):

you are welcome